a line has a slop of 4/5 and passes through (1,3). What is the slope-intercept form of the equation of this line?
Answer by lwsshak3(11628)

a line has a slop of 4/5 and passes through (1,3). What is the slope-intercept form of the equation of this line?
**
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
equation of given line:
y=4x/5+b
solve for b using coordinates of given point(1,3) on the line
3=4*1/5+b
15=4+5b
5b=11
b=11/5
Equation of given line:
y=4x/5+11/5